karina was comparing the volume of a 20-inch cube to the volume of a 10-inch cube . how many times greater is the volume of the larger cube?

Answer by TimothyLamb(4379)

x = volume of 20 inch cube
y = volume of 10 inch cube
---
x = 20*20*20
y = 10*10*10
---
x/y = (20*20*20)/(10*10*10)
---
answer:
x/y = 8 times
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php